OpenStudy (anonymous):

Have an algebra question: solve by completing the square: x^2-4x+11 = 0 Can anyone help me with the answer?

OpenStudy (anonymous):

yes lets to it the easy way

OpenStudy (anonymous):

i'll check it at the end for you satellite73

OpenStudy (anonymous):

\[x^2-4x=-11\] \[(x-2)^2=-11+2^2=-11+4=-7\]

OpenStudy (anonymous):

now if you are working with real numbers you can stop right there, because the left hand side is a prefect square and the right hand side is negativef

OpenStudy (anonymous):

if you are working with comlex numbers you write \[(x-2)^2=-7\] \[x-2=\pm\sqrt{-7}\] \[x-2=\pm\sqrt{7}i\] \[x=2\pm\sqrt{7}i\]

OpenStudy (anonymous):

so basically pw8589, what he's doing is isolating the x terms on one side and the constants on the other side. then divide you single cx (c standing for a constant) by 2 and square that number

OpenStudy (anonymous):

I followed as x^2 -4x = -11 Then went to (x-2)^2 = -9 SR (x-2)^2 = +/- sr of -9 am I on the right track?

OpenStudy (anonymous):

(x-2)^2 = -9 is not correct... you added 4 on the one side so you must add 4 on the other side.. you have -9, it should be -7

OpenStudy (anonymous):

When taught completing the square - we divided the 4x by 2... which is where I was getting the 2 from instead of the 4

OpenStudy (anonymous):

look at it this way \[(x^2-4x)+4=-11 + 4\]

OpenStudy (anonymous):

Thank you